'y' = -2/9x
find 'x' when 'y' = -34

Answer by checkley77(12569)

'y' = -2/9x
find 'x' when 'y' = -34
-34=-2/9X
X=-34/(-2/9)
X=-34*-9/2
X=17*9
X=153 ANS.